The following data relates to the market value of economic transactions for the three main sectors of your country’s economy.

SECTOR VALUE OF OUTPUT PURCHASES FROM OTHER FIRMS
Agriculture 300 160
Manufacturing 200 150
Services 150 120

Required:

(i) Compute the gross domestic product (GDP) and explain the method used and why.

(ii) Given that: depreciation = 90, indirect taxes = 70, subsidies = 30, payments to factors of production from abroad = 20, payments to foreign factors = 40.

Compute:

1) Gross national product (GNP) at market prices.

2) Net national product (NNP) at market prices.

3) Net national product (NNP) at factor cost.

4) Net domestic product (NDP) at factor cost
1
Expert's answer
2020-06-29T16:40:49-0400

I) GDP

Income Method - Adding all factor payments.

GDP=Value of Output +Purchases

=(300+200+120)+(160+150+120)

650+430=1080.

GDP =1080.

2i)GNP(at Market Prices) =GDP+Net factor Income from Abroad.

=1080-(40-20).

=1100

ii)NNP(at Market Prices) =GDP - Depreciation.

=1080-90

=990

iii)NNP (at Factor Cost)

=NNP(at MP) - Indirect Taxes

=990-70

=920

iv)NDP(at Factor Cost) =NNP (FC) +Subsidies

=920 +30

=950.